OpenStudy (mtepecz):

Solve. cot^2x+3cotx=4

OpenStudy (dayakar):

cot^2x+3cotx-4=0 let cot x = a it becomes a^2+3x-4=0 it is like quadratic equation can u solve for a

OpenStudy (dayakar):

sorry i typed x a^2+3a-4=0